OpenStudy (anonymous):

Question 4 (Multiple Choice Worth 1 points) The graph of a system of parallel lines will have no solutions. Always Sometimes Never

OpenStudy (anonymous):

Always

OpenStudy (anonymous):

CAN I ASK YOU ONE MORE ??

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

The graph of a system of equations with the same slope will have no solutions. Always Sometimes Never

OpenStudy (anonymous):

always

OpenStudy (anonymous):

TY

OpenStudy (anonymous):

AND ONE MORE 2x + 3y = 3 3x – 2y = 11 What is the solution to the system? (–3, 3) (–1, –7) (1, –4) (3, –1)

OpenStudy (anonymous):

multiply first eqn by 3 and second by -2 6x + 9y = 9 -6x + 4y = -22 now add the two 13y = -13 y = -13/13 = -1 substitute y = -1 in first eqn 2x + 3(-1) = 3 2x - 3 = 3 2x = 6 x = 6/2 = 3 so solution is (3, -1)

OpenStudy (anonymous):

both say "will have no solution"

OpenStudy (radar):

"never mind" lol, I see now, good answer Harkirat

OpenStudy (anonymous):

THANK YOU THATS WHAT I HAD JUST HAD TO MAKE SURE CAN YOU HELP ME WITH JUST ONE MROE QUESTION THEN ILL STOP PBOTHER YOU

OpenStudy (radar):

Go ahead, I will try.

OpenStudy (radar):

It is no bother.

OpenStudy (anonymous):

Part 1: Provide a system of TWO equations in slope-intercept form, with only one solution. Using complete sentences, explain why this system has one solution. Part 2: Provide a system of TWO equations in slope-intercept form with no solutions. Using complete sentences, explain why this system has no solutions. Part 3: Provide a system of TWO equations in slope-intercept form with infinitely many solutions. Using complete sentences, explain why this system has infinitely many solutions. ** ALL I NEEED IS EXAMPLES

OpenStudy (radar):

O.K. I'm on it.

OpenStudy (anonymous):

OK THANK YOU

OpenStudy (anonymous):

I GOT A ANSWER TO IT I WANT YOU TO TELL ME IS IT RIGHT

OpenStudy (anonymous):

for Part 1, system with one solution, When you graph out a system of equations ( let say there is only 2). The intercept of the 2 lines, the point where they intercept is the solution. So if they intercept at only one point there is only one solution. EXP., y=x+2 y=2x+4.. For part 2, system with no solution, for 2 lines to not intersect, is when both lines are parallel. if both lines are parallel they will never intersect even when stretched to infinity. EXP.y=x+2 y=x+4 for part 3, system with infinite solution, it is when both lines intersect at every point, which means to say they are the same line An example will be x+y=2 2x+2y=4 Both are the same line when simplified. EXP y=x+1 2y=2x+2

OpenStudy (radar):

Pt. 1:y=3x+1 y=4x-1 Pt. 2. y=-2x +9 y= -2x + 2 pt. 3 y = x -1 y= (2) .5x +(-1)

OpenStudy (radar):

Normally when an equation is in slope intercept form, the coefficient of y is a 1. So for Pt. 3, it is the same equation (an identity) and has infinite solutions.

OpenStudy (anonymous):

if you dont mind i have more quest i reallly appericate your help